A four-way set-associative cache has lines of 32 bytes and a total size of 4 kB. The 32-MB main memory is byte addressable. Show the format of main memory addresses.
Answer #1

Number of block at the cache is = 4KB /32 BYTES= 4 * 1024 / 32 = 128 BLOCK

SET LINE = 128 /4 = 32 SET OF 4 LINE EACH

32 = 2^5 = 5 BIT FOR SET

MAIN MEMEORY = 32MB = 2^25 = 25 BIT

NUMBER OF BLOCK AT MEMORY = 32 MB / 32 BYTES= 2^20 = 20 BITS

TAG FILED= 20- 5=15

BLOCK SIZE = 32 BYTE= 5 BIT FOR WORD

SO MEMEORY FORMAT

TAG -> 15 BITS

SET -> 5

WORD(W)->5